OpenStudy (anonymous):

please helpme!! The first-, second-, and third-year enrollment values for a technical school are shown in the table below. Enrollment at a Technical School Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table?

OpenStudy (anonymous):

the graphs cae out warped but the columns are supposed to be under each year 2nd and 3rd column go under year 2 and 3 answer choices are The solution to f(x) = s(x) is x = 2,009. The solution to f(x) = s(x) is x = 785. The solution to f(x) = t(x) is x = 2,010. The solution to f(x) = t(x) is x = 740.

OpenStudy (anonymous):

@Hero

OpenStudy (sohailiftikhar):

it's c

OpenStudy (anonymous):

can you explain please i really hate taking answers because i usually get multiple of the same type of questions

OpenStudy (sohailiftikhar):

you can see from data when x=2010 f(x)=740=t(x)

OpenStudy (anonymous):

ok thank you

OpenStudy (sohailiftikhar):

np